# Weak measurement-based protocol for ergotropy protection in open quantum batteries

### v1

@@ Abstract
Quantum batteries are emerging as highly efficient energy storage devices that can exceed classical performance limits. Although there have been significant advancements in controlling these systems, challenges remain in stabilizing stored energy and minimizing losses due to inevitable environmental interaction. In this paper, we propose a protocol that employs selective weak measurements to protect quantum states from such influence and mitigate battery discharging. We establish thermodynamic constraints that allow this method to be implemented without disrupting the overall energy and ergotropy balance of the system. Our findings demonstrate that appropriately chosen measurement intensity can reduce unwanted discharging effects, thereby preserving ergotropy and improving the stability of quantum batteries. Additionally, we explore how weak measurements influence the coherent and incoherent components of ergotropy, providing new insights into the practical application of quantum coherence in energy storage technologies.
@@ I Introduction
In recent years, quantum batteries have emerged as promising devices for energy storage. These systems, composed of quantum states as energy carriers, aim to employ genuine quantum properties, such as coherence and entanglement, to surpass classical limits [1, 2, 3, 4, 5, 6]. In this sense, quantum batteries were shown to exhibit unique features such as coherence-driven enhancements in ergotropy and charging processes [6, 7, 2, 8]. Significant progress has been made in understanding and manipulating these devices, with studies exploring diverse approaches like collisional models [9, 10], super- and sub-radiance [11], reservoir engineering [12, 13], phase transitions [14], micro-maser quantum batteries [15, 16], topological features [17, 18], catalysis [19], non-Markovian effects [20], optimal control strategies [21, 22, 23], and charging by measurement [24].
@@
Various physical platforms have been proposed as viable candidates for quantum energy storage. Under cryogenic conditions, technologies such as superconducting circuits [25, 26, 27, 23], semiconductor quantum dots [28], ultracold-atoms [29], and nitrogen-vacancy centers in diamond [30] provide promising options, enabling quantum coherence and collective behavior necessary for efficient charging. On the other hand, room-temperature approaches, including organic micro-cavities [5] and nuclear spin systems [31, 32], showcase the potential for robust quantum energy storage through mechanisms that minimize decoherence, marking key advancements toward practical quantum batteries across a range of operating conditions.
@@
Nevertheless, despite ongoing efforts, significant challenges persist, particularly in more realistic scenarios where quantum coherence is susceptible to decoherence effects [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. In this context, a charged quantum battery exposed to a dissipative environment renders not only the irreversible leakage of energy from the battery into the environment but also the loss of its charge, commonly quantified by the ergotropy, i.e., the extractable work from a quantum state [44]. For instance, the contact with a thermal reservoir leads the battery to a Gibbs state, which is completely passive [45]. Such effects are particularly detrimental to the performance of quantum technologies, as they limit the ability to harness ergotropy for practical applications. Therefore, characterizing and mitigating the influence of thermal reservoirs is essential for the reliable operation of quantum devices in general [46, 47, 48]. As proposed by [49], a possible way to preserve a quantum system from decaying is given by the continuous observation of the quantum system by a sequence of frequent projective measurements (a feature of the Zeno effect [50]), which can hinder the discharging process, effectively freezing the energy stored in a quantum battery. However, this method may not be experimentally friendly due to the required frequent state observation.
@@
Furthermore, measurements are also being incorporated in quantum thermodynamics, e.g., as additional strokes in thermal cycles, with measurement-powered engines exploiting the fact that the measurement action generally disturbs the observed system, thereby altering its internal energy [51, 52, 53]. By employing non-selective measurements, for instance, a single temperature heat engine without feedback control to extract work was introduced [51], and several other measurement-based protocols have been explored  [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. In particular, since generalized measurements can be implemented as a sequence of weak measurements [80, 81], their intensity can be continuously adjusted to transition between weak and strong measurement regimes [82, 83]. Thus, weakly measured systems were shown to be crucial for consistently observing work and heat contributions in an externally driven quantum stochastic evolution [84] and were employed to develop a deterministic protocol based on non-selective variable strength measurements to perform a heat engine cycle [65, 78, 85]. Moreover, considering a bipartite correlated quantum system, selective weak measurements can offer more work extraction than the one described by the daemonic ergotropy  [86], i.e., the amount of extractable work that arises from information gained in a strong (projective) measurement [87].
Here, we investigate the potentiality of weak measurements in attenuating environment-induced losses, aiming to harness quantum mechanical subtleties for more efficient energy storage and extraction in two-level systems-based quantum batteries. The protocol builds on the structure of previous theoretical and experimental studies on protecting quantum states from decoherence using weak measurements and measurement reversal operations [88, 89, 90, 91, 92, 93, 94, 95]. The key distinction lies in our incorporation of thermodynamic-based constraints, enabling these methods to be applied without altering the overall energetics of the protocol. Specifically, before the discharging process, an initial weak measurement is performed, modifying the system energy and ergotropy. Subsequently, the charged system undergoes dissipative dynamics, during which a second weak measurement is carried out to compensate for the previously induced energetic and ergotropic changes. Despite the net exchange of these quantities being null, a positive gain can be observed at the end of the protocol. The feasibility of such a procedure is a direct consequence of the symmetry of the weak measurement and its reversal. Notably, without affecting the protocol’s overall probability of success, quantum coherence can be employed to enhance the measurement-based battery recovery process from discharging effects, as demonstrated in this work.
The remainder of this article is structured as follows. In Sec. II we introduce the physical context and model to explain the dissipation loss mitigation protocol. In Sec. III we present our main results by analyzing the protocol performance for the incoherent and coherent part of the ergotropy, as well as the total ergotropy, and identify parameter regions where the protocol yields promising outcomes. Finally, in Sec. IV, we summarize our main conclusions and discuss potential perspectives for future work.
II Ergotropy loss mitigation protocol
@@
Figure 1: Ergotropy-protecting protocol based on weak measurements: We consider a qubit quantum battery initially charged, which then undergoes a discharging process. Before this, a weak measurement with strength $m$ is applied in the first step, changing the system’s ergotropy (charge) and energy (dashed blue line). In the intermediate step, during the discharging phase, a second reversal weak measurement with strength $w$ is performed, compensating for the previous changes. Despite that, by carefully tuning the measurement intensities, the protocol can offer an advantage over a purely discharging process (continuous red line).
@@
$\mathcal{R}[\hat{\rho}]\coloneqq\textrm{Tr}\left[\left(\hat{\rho}-\hat{\rho}_{p}\right)\hat{H}\right]=\sum_{k,j}^{d}E_{k}p_{j}\left(|\langle E_{k}|j\rangle|^{2}-\delta_{k,j}\right),$ (1)
where $\hat{\rho}_{p}=\sum_{i}^{d}p_{i}|E_{i}\rangle\langle E_{i}|$ is the passive state, which is unitarily related to the initial state $\hat{\rho}$. Generally, the battery state may exhibit coherences in the $\hat{H}$ basis, which is a unique quantum resource [96]. Identifying the contribution of coherence to the battery ergotropy is therefore insightful. To capture this, the coherent ergotropy within $\hat{\rho}$ is defined as [97, 98, 99]
@@
where $\mathcal{R}^{inc}[\hat{\rho}]$ represents the incoherent component of the ergotropy, such that $\mathcal{R}^{inc}[\hat{\rho}]=\mathcal{R}[\hat{\rho}^{diag}]$, where $\hat{\rho}^{diag}=\sum_{k}^{d}\langle E_{k}|\hat{\rho}|E_{k}\rangle|E_{k}\rangle\langle E_{k}|$ denotes the diagonal (incoherent) state of $\hat{\rho}$, obtained by removing its off-diagonal elements. Note that, by definition, $\mathcal{R}^{coh}[\hat{\rho}]$ is zero for incoherent states.
II.1 Setup and protocol description
@@
Notice that $\rho_{m}(0)=\rho_{0}$ $\forall m$ for a fully charged battery, i.e., $P_{0}=1$ and $Q_{0}=0$. In this particular case, the action of $\hat{M}_{m}$ has no effect. However, due to the non-unitary nature of thermalization dynamics, we can generally have $P_{0}<1$ and proceed with the protocol; (iii) During a dissipation time $t\in[0,\tau]$, the battery is assumed to interact with the external environment. In this interval, its dynamics are governed by Eq. (3). Consequently, the state of the battery evolves as
@@
It is important to emphasize that the measurements are assumed to be instantaneous. Also, note that the diagonal states and probabilities at each step are independent of the initial coherence $Q_{0}$. This independence arises from the commutation relations $[\hat{M}_{m},\hat{H}]=[\hat{W}_{w},\hat{H}]=0$ and from Eq. (3), implying that the population changes induced by the successive actions of $\hat{M}_{m}$, thermalization dynamics, and $\hat{W}_{w}$ do not couple them with the off-diagonal elements. The opposite, however, does not hold; that is, the coherence is coupled with the population during the measurements.
@@
is the threshold strength for obtaining a finite final incoherent ergotropy, i.e., one has to guarantee $w>w^{\prime}$ to have an active post-reversal measurement diagonal state. Additionally, the coherent component can be written as
$\mathcal{R}_{mw}^{coh}(\tau)=\frac{1}{2}\left(\psi_{mw}(\tau)-\sqrt{\psi_{mw}^{2}(\tau)-4\left|Q_{mw}(\tau)\right|^{2}}\right),$ (15)
where $\psi_{mw}(\tau)\coloneqq\sqrt{2\mu[\hat{\rho}_{mw}(\tau)]-1}$, with $\mu[\hat{\rho}]\coloneqq\textrm{Tr}\left[\hat{\rho}^{2}\right]$ being the purity of the state $\hat{\rho}$.
@@ II.2 Protocol working regime
Note that the weak measurement and its reversal generally disturb the battery state; specifically, the former brings the battery closer to the ground state, while the latter pushes it toward the excited state. As a result, these protocol steps inevitably induce energy changes, potentially charging (or discharging) the battery. We therefore define $\varepsilon_{mw}(\tau)$ as the energy shift of the battery due to the measurements as
@@
$\begin{split}\Delta E_{m}&\coloneqq\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{m}(0)-\hat{\rho}_{0}\right)\right],\\ \Delta E_{mw}(\tau)&\coloneqq\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{mw}(\tau)-\hat{\rho}_{m}(\tau)\right)\right].\end{split}$ (20)
It is worth highlighting that such quantity is independent of the initial coherence of the battery $Q_{0}$. In fact, $\varepsilon_{mw}(\tau)$ can be defined and computed only in terms of the diagonal (incoherent) component of the states. These changes may also be accompanied by an ergotropy shift $\mathcal{W}_{mw}(\tau)$, such that
@@
$\begin{split}\Delta\mathcal{R}_{m}&\coloneqq\mathcal{R}[\hat{\rho}_{m}(0)]-\mathcal{R}[\hat{\rho}_{0}],\\ \Delta\mathcal{R}_{mw}(\tau)&\coloneqq\mathcal{R}[\hat{\rho}_{mw}(\tau)]-\mathcal{R}[\hat{\rho}_{m}(\tau)].\end{split}$ (22)
Hence, to ensure that the protocol does not actively provide a net external resource to the battery and effectively (dis)charge it, one should specify the measurement strengths $m$, $w$, and the dissipation time $\tau$ such that
$\varepsilon_{mw}(\tau)=0,\qquad\mathcal{W}_{mw}(\tau)=0.$ (23)
The simultaneous validation of both requirements restricts the parameter space to valid working regions for our proposed protocol. In particular, for fixed values of $\gamma$, $f$, $P_{0}$, $m$, and $\tau$, one can show that the reversal measurement strength $\tilde{w}$ needed to restore the energy lost due to the weak measurement is given by
$\tilde{w}=\textrm{N}_{m}\frac{1-P_{0}-\textrm{N}_{m}\left(1-f\right)+e^{\gamma\tau}\left[\textrm{N}_{m}\left(P_{m}(\tau)+P_{0}-1-f\right)+1-P_{0}\right]}{\left[\textrm{N}_{m}(P_{m}(\tau)+P_{0}-1)+1-P_{0}\right]\left[\textrm{N}_{m}(1-f)\left(e^{\gamma\tau}-1\right)+1-P_{0}\right]},$ (24)
which ensures null energetic shift, $\varepsilon_{m\tilde{w}}(\tau)=0$. Also, one should restrain into physical regions, characterized by strengths satisfying $\tilde{w}\in[0,1]$. Hence, given Eq.  (24), one can also find the subspace with null ergotropic shift, satisfying $\mathcal{W}_{m\tilde{w}}(\tau)=0$ (see Appendix B for further discussion).
@@
In short, a successful and suitable application of the described protocol is characterized both by (i) a positive ergotropy gain, ¯$\overline{\mathcal{R}}_{mw}(\tau)>0$, and (ii) zero energy and ergotropy shifts, $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=0$. In the following, we present our main results, demonstrating that this protocol yields an ergotropy gain, effectively mitigating battery discharging when the two measurement intensities and dissipation time are appropriately selected. Additionally, we identify the impact of both the coherent and incoherent components of the battery ergotropy. It will be assumed $\omega=1$, $f=0.3$ and $\gamma=10^{-2}$ for the following analyzes.

### v2

# Two-time weak measurement protocol for ergotropy protection in open quantum batteries

@@ Abstract
Quantum batteries are emerging as highly efficient energy storage devices that can exceed classical performance limits. Although there have been significant advancements in controlling these systems, challenges remain in stabilizing stored energy and minimizing losses due to inevitable environmental interaction. In this paper, we propose a protocol that employs selective weak measurements to protect quantum states from such influence and mitigate battery discharging, that is feasible in state-of-the-art technologies. We establish thermodynamic constraints that allow this method to be implemented without disrupting the overall energy and ergotropy balance of the system, i.e., with no extra net recharging. Our findings demonstrate that appropriately chosen measurement intensity can reduce unwanted discharging effects, thereby preserving ergotropy and improving the stability of quantum batteries. We illustrate the protocol with single and two-qubit systems and establish the generalization for $N$-cell batteries. Additionally, we explore how weak measurements influence the coherent and incoherent components of ergotropy, providing new insights into the practical application of quantum coherence in energy storage technologies.
@@ I Introduction
In recent years, quantum batteries have emerged as promising devices for energy storage. These systems, composed of quantum states as energy carriers, aim to employ genuine quantum properties, such as coherence and entanglement, to surpass classical limits [1, 2, 3, 4, 5, 6, 7]. In this sense, quantum batteries were shown to exhibit unique features such as coherence-driven and collective enhancements in ergotropy and charging processes [5, 8, 2, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Significant progress has been made in understanding and manipulating these devices, with studies exploring diverse models [19, 20, 21, 22, 23, 14, 24, 25] and approaches like collisional models [26, 27], super- and sub-radiance [28], reservoir engineering [29, 30], phase transitions [31], micro-maser quantum batteries [32, 33], topological features [34, 35], catalysis [36], non-Markovian effects [37], wireless charging [38, 39, 40], reinforcement learning [41, 42], charging by measurement [43], and optimal control strategies [44, 45, 46].
@@
Various physical platforms have been proposed as viable candidates for quantum energy storage. Under cryogenic conditions, technologies such as superconducting circuits [47, 48, 49, 46, 50, 51, 52], semiconductor quantum dots [53] and ultracold-atoms [54] provide promising options, enabling quantum coherence and collective behavior necessary for efficient charging. On the other hand, room-temperature approaches, including organic micro-cavities [4], optical setups [55], nitrogen-vacancy (NV) centers in diamond [56] and nuclear spin systems [57, 58], showcase the potential for robust quantum energy storage through mechanisms that minimize decoherence, marking key advancements toward practical quantum batteries across a range of operating conditions.
@@
Nevertheless, despite ongoing efforts, significant challenges persist, particularly in more realistic scenarios where quantum coherence is susceptible to decoherence effects [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 38, 71, 72]. In this context, a charged quantum battery exposed to a dissipative environment renders not only the irreversible leakage of energy from the battery into the environment but also the loss of its charge, commonly quantified by the ergotropy, i.e., the extractable work from a quantum state [73]. For instance, the weak contact with a thermal reservoir leads the battery to a Gibbs state, which is completely passive [74]. Such effects are particularly detrimental to the performance of quantum technologies, as they limit the ability to harness ergotropy for practical applications. Therefore, characterizing and mitigating the influence of thermal reservoirs is essential for the reliable operation of quantum devices in general [75, 76, 77, 78, 79]. As proposed by [80], a possible way to preserve a quantum system from decaying is given by the continuous observation of the quantum system by a sequence of frequent projective measurements (a feature of the Zeno effect [81]), which can hinder the discharging process, effectively freezing the energy stored in a quantum battery. However, this method may not be experimentally friendly due to the required frequent state observation.
@@
Furthermore, measurements are also being incorporated in quantum thermodynamics, e.g., as additional strokes in thermal cycles, with measurement-powered engines exploiting the fact that the measurement action generally disturbs the observed system, thereby altering its internal energy [82, 83, 84]. For instance, by employing non-selective measurements, a single temperature heat engine without feedback control to extract work was introduced [82], and several other measurement-based protocols have been explored  [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110]. In particular, since generalized measurements can be implemented as a sequence of weak measurements [111, 112], their intensity can be continuously adjusted to transition between weak and strong measurement regimes [113, 114]. Thus, weakly measured systems were shown to be crucial for consistently observing work and heat contributions in an externally driven quantum stochastic evolution [115] and were employed to develop a deterministic protocol based on non-selective variable strength measurements to perform a heat engine cycle [96, 109, 116]. Considering bipartite quantum systems, local strong (projective) measurements were shown to improve the quantum battery capacity [117], while selective weak measurements can offer more work extraction than the one described by the daemonic ergotropy  [118], i.e., the amount of extractable work that arises from information gained in a strong measurement [119]. In this sense, weak measurements have been widely adopted across diverse quantum platforms, with successful implementations in nuclear magnetic resonance (NMR) [120, 109, 121], NV centers [122], trapped ion systems [113], photonic systems [123, 124, 125], and superconducting qubits [126, 127, 128, 129, 130], demonstrating their versatility for contemporary quantum applications. The thermodynamic cost of implementing weak measurements fundamentally depends on the underlying protocol and the physical nature of the probe, necessitating an analysis of interconnected factors such as measurement strength, energy consumption, information extraction, and thermodynamic irreversibility [131, 132]. Furthermore, renormalization procedures used in post-measurement state analysis introduce additional energetic considerations, which are essential to characterize the impact of measurements on the performance of such thermodynamic protocols.
@@
Figure 1: Two-time Weak Measurement (TWM) protocol: We consider a quantum battery initially charged, which then undergoes a discharging process. Before this, a weak measurement with strength $m$ is applied in the first step, changing the system’s energy and ergotropy (charge), represented by the dashed blue line. In the intermediate step, during the discharging phase, a second reversal weak measurement with strength $w$ is performed at instant $\tau$, compensating for the previous changes. By carefully tuning the measurement intensities, the protocol can offer an advantage despite null net recharging, over a purely discharging process undergone by $\mathcal{R}[\hat{\rho}(t)]$ (continuous red line), which is quantified by the ergotropy gain ¯$\overline{\mathcal{R}}_{mw}(\tau)\coloneqq\mathcal{R}_{mw}(\tau)-\mathcal{R}[\hat{\rho}(\tau)]$.
Here, we investigate the potentiality of weak measurements in attenuating environment-induced losses, aiming to harness quantum mechanical subtleties for more efficient energy storage and extraction in two-level systems-based quantum batteries. The two-time weak measurement (TWM) protocol builds on the structure of previous theoretical and experimental studies on protecting quantum states from decoherence using weak measurements and measurement reversal operations [124, 133, 125, 134, 135, 136, 137, 138, 139]. The key distinction lies in our incorporation of thermodynamic-based constraints applied on the target system while considering the measurements as resources, enabling these methods to be applied without altering the overall energetics of the protocol. Specifically, as shown in Fig. 1, before the discharging process, an initial weak measurement is performed, modifying the system energy and ergotropy. Subsequently, the charged system undergoes dissipative dynamics, during which a second weak measurement is carried out to compensate for the previously induced energetic and ergotropic changes. Despite the net exchange of these quantities being constrained to be null, a positive gain can be observed at the end of the protocol by tuning the relevant parameters. The feasibility of such a procedure is a direct consequence of the symmetry of weak measurement and its reversal, and the different discharging rates resembling the ergotropic counterpart of the Mpemba effect [140, 141]. We separated the analysis to consider both the isolated incoherent and coherent components of ergotropy. Notably, without affecting the protocol’s overall probability of success, quantum coherence can be employed to enhance the measurement-based battery recovery process from discharging effects, as suggested in this work. The protocol is investigated for batteries composed of single and two-qubit systems, and the generalization for $N$-cells batteries is introduced.
The remainder of this article is structured as follows. In Sec. II we introduce the physical context and model to explain the dissipation loss mitigation protocol. In Sec. III we present our main results by analyzing the protocol performance for the incoherent and coherent part of the ergotropy, as well as the total ergotropy, and identify parameter regions where the protocol yields promising outcomes. In Sec. IV we describe how the protocol is extended to multi-qubit quantum batteries, and illustrate the procedure for a two-qubit system with quantum correlations. Finally, in Sec. V, we summarize our main conclusions and discuss potential perspectives for future work.
II Ergotropy loss mitigation: The Two-time Weak Measurement (TWM) protocol
@@
$\mathcal{R}[\hat{\rho}]\coloneqq\textrm{Tr}\left[\left(\hat{\rho}-\hat{\rho}^{p}\right)\hat{H}\right]=\sum_{k,j}^{d}E_{k}p_{j}\left(|\langle E_{k}|j\rangle|^{2}-\delta_{k,j}\right),$ (1)
where $\hat{\rho}^{p}=\sum_{i}^{d}p_{i}|E_{i}\rangle\langle E_{i}|$ is the passive state, which is unitarily related to the initial state $\hat{\rho}$. Generally, the battery state may exhibit coherences in the $\hat{H}$ basis, which is a unique quantum resource [142]. Identifying the contribution of coherence to the battery ergotropy is therefore insightful. To capture this, the coherent ergotropy within $\hat{\rho}$ is defined as [143, 144, 145]
@@
where $\mathcal{R}^{inc}[\hat{\rho}]$ represents the incoherent component of the ergotropy, such that $\mathcal{R}^{inc}[\hat{\rho}]=\mathcal{R}[\hat{\rho}^{diag}]$, where $\hat{\rho}^{diag}=\sum_{k}^{d}\langle E_{k}|\hat{\rho}|E_{k}\rangle|E_{k}\rangle\langle E_{k}|$ denotes the diagonal (incoherent) state of $\hat{\rho}$, obtained by removing its off-diagonal elements in the energy basis. Note that, by definition, $\mathcal{R}^{coh}[\hat{\rho}]$ is zero for incoherent states.
II.1 Setup and the TWM protocol description
@@
Notice that $\rho_{m}(0)=\rho_{0}$ $\forall m$ for a fully charged battery, i.e., $P_{0}=1$ and $Q_{0}=0$. In this particular case, the action of $\hat{M}_{m}$ has no effect. However, due to the non-unitary nature of thermalization dynamics, one can wait for $P_{0}<1$ and proceed with the protocol; (iii) During a dissipation time $t\in[0,\tau]$, the battery is assumed to interact with the external environment. In this interval, its dynamics are governed by Eq. (3). Consequently, the state of the battery evolves as
@@
It is important to emphasize that the measurements are assumed to be instantaneous. While measurements inherently require finite time, our previous assumption holds valid when the system’s characteristic dynamic timescale exceeds the measurement duration. Experimental evidence from both NMR [109] and superconducting qubits [126, 127, 128, 130] corroborates that weak measurements can be performed in time-frames significantly shorter than the relevant relaxation and decoherence times. Also, note that the diagonal states and probabilities at each step are independent of the initial coherence $Q_{0}$. This independence arises from the commutation relations $[\hat{M}_{m},\hat{H}]=[\hat{W}_{w},\hat{H}]=0$ and from Eq. (3), implying that the population changes induced by the successive actions of $\hat{M}_{m}$, thermalization dynamics, and $\hat{W}_{w}$ do not couple them with the off-diagonal elements. The opposite, however, does not hold; that is, the coherence is coupled with the population during the measurements.
@@
is the threshold strength for obtaining a finite final incoherent ergotropy, i.e., one has to guarantee $w>w^{\prime}$ to have an active post-reversal measurement diagonal state. In general, the coherent component for two-level systems with Hamiltonian $\hat{H}=\omega|e\rangle\langle e|$ can be written as
$\mathcal{R}^{coh}[\hat{\rho}]=\frac{\omega}{2}\left(\psi-\sqrt{\psi^{2}-4\left|Q\right|^{2}}\right),$ (15)
where $\psi\coloneqq\sqrt{2\mu[\hat{\rho}]-1}$, with $\mu[\hat{\rho}]\coloneqq\textrm{Tr}\left[\hat{\rho}^{2}\right]$ being the purity of the state $\hat{\rho}$, and $Q$ is the off-diagonal element. Thus, at the end of the protocol, one has $\mathcal{R}_{mw}^{coh}(\tau)=\mathcal{R}^{coh}[\hat{\rho}_{mw}(\tau)]$.
@@ II.2 Protocol working regime
Note that the weak measurement and its reversal generally disturb the battery state; specifically, the former brings the battery closer to the ground state, while the latter pushes it toward the excited state. As a result, these protocol steps inevitably induce energy changes, potentially charging (or discharging) the battery. We therefore define $\varepsilon_{mw}(\tau)$ as the energy shift of the battery, quantifying the total (net) energy change induced by the measurements, i.e., as
@@
$\begin{split}\Delta E_{m}&\coloneqq\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{m}(0)-\hat{\rho}_{0}\right)\right],\\ \Delta E_{mw}(\tau)&\coloneqq\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{mw}(\tau)-\hat{\rho}_{m}(\tau)\right)\right],\end{split}$ (20)
are the energy changes induced by the weak and reversal measurements, respectively. It is worth highlighting that this quantity is independent of the initial coherence of the battery $Q_{0}$. In fact, $\varepsilon_{mw}(\tau)$ can be defined and computed only in terms of the diagonal (incoherent) component of the states.
These energetic changes are expected to be accompanied by ergotropic counterparts. Thus, to account for the total measurement-induced ergotropy difference, we define the ergotropy shift $\mathcal{W}_{mw}(\tau)$ as
@@
$\begin{split}\Delta\mathcal{R}_{m}&\coloneqq\mathcal{R}[\hat{\rho}_{m}(0)]-\mathcal{R}[\hat{\rho}_{0}],\\ \Delta\mathcal{R}_{mw}(\tau)&\coloneqq\mathcal{R}[\hat{\rho}_{mw}(\tau)]-\mathcal{R}[\hat{\rho}_{m}(\tau)],\end{split}$ (22)
are the individual ergotropy contributions for each measurement. Alternatively, from the ergotropy definition given by Eq. (1), one can rewrite the ergotropy shift as $\mathcal{W}_{mw}(\tau)=\varepsilon_{mw}(\tau)-\varepsilon_{mw}^{p}(\tau)$, where $\varepsilon_{mw}^{p}(\tau)$ represents the total energetic changes relative to the passive states of each step, i.e. $\varepsilon_{mw}^{p}(\tau)=\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{m}^{p}-\hat{\rho}_{0}^{p}\right)\right]+\textrm{Tr}\left[\hat{H}\left(\hat{\rho}_{mw}^{p}(\tau)-\hat{\rho}_{m}^{p}(\tau)\right)\right]$.
Observe that if at the end of the protocol one has induced net positive (negative) ergotropy, part of the recovered (loss) charge is due to the measurements themselves, which means that the procedure would be directly (dis)charging the battery. Hence, to ensure that the TWM protocol does not actively provide a net external resource to the battery and effectively (dis)charge it, one should specify the measurement strengths $m$, $w$, and the dissipation time $\tau$ such that
$\begin{split}\varepsilon_{mw}(\tau)&=0,\\ \qquad\mathcal{W}_{mw}(\tau)&=-\varepsilon_{mw}^{p}(\tau)=0.\end{split}$ (23)
The simultaneous validation of both requirements restricts the parameter space to valid working regions for our proposed protocol. In fact, given Eq. (20) for the energetic shift, and fixed values of $\gamma$, $f$, $P_{0}$, weak measurement strength $m$, and dissipation time $\tau$, one can find the expression of the reversal measurement strength $\tilde{w}$ needed to restore the energy lost due to the weak measurement, i.e., from the constraint for $\varepsilon_{m\tilde{w}}(\tau)=0$, one can invert the relation and show that
$\tilde{w}=\textrm{N}_{m}\frac{1-P_{0}-\textrm{N}_{m}\left(1-f\right)+e^{\gamma\tau}\left[\textrm{N}_{m}\left(P_{m}(\tau)+P_{0}-1-f\right)+1-P_{0}\right]}{\left[\textrm{N}_{m}(P_{m}(\tau)+P_{0}-1)+1-P_{0}\right]\left[\textrm{N}_{m}(1-f)\left(e^{\gamma\tau}-1\right)+1-P_{0}\right]}$ (24)
ensures a null energetic shift. In addition, we should limit ourselves to physical regions, characterized by strengths that satisfy $\tilde{w}\in[0,1]$. Hence, given Eq.  (24), one can also find the subspace with null ergotropic shift, satisfying $\mathcal{W}_{m\tilde{w}}(\tau)=0$ (see Appendix B for further discussion).
@@
Figure 2: Incoherent ergotropy gain as a function of the weak measurement strength $m$ and initial population $P_{0}$ along the null energetic shift reversal measurement strength $\tilde{w}\in[0,1]$ for distinct dissipation times (a) $\tau=\tau_{\gamma}/2$, (b) $\tau=\tau_{\gamma}$, and (c) $\tau=10\tau_{\gamma}$. The red-hatched region and red curves indicate that $\mathcal{W}_{m\tilde{w}}(\tau)=0$. The $\eta_{2}$ lines are highlighted.
@@
In short, a successful and suitable application of the described TWM protocol is characterized both by (i) a positive ergotropy gain, ¯$\overline{\mathcal{R}}_{mw}(\tau)>0$, and (ii) zero energy and ergotropy shifts, $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=0$. While the former means the TWM protocol provided some advantage, the latter guarantees this gain is not due to extra net recharging. In the following, we present our main results, demonstrating that this protocol yields an ergotropy gain, effectively mitigating battery discharging when the two measurement intensities and dissipation time are appropriately selected. Additionally, we identify the impact of both the coherent and incoherent components of the battery ergotropy. Without loss of generality and unless stated otherwise, it will be assumed $f=0.3$ and $\gamma=10^{-2}$ for the following analyzes.